OpenStudy (anonymous):

Find the average rate of change of f(x)=x^3-3x+1 (a) From -5 to -3 (b) From -2 to 2 (c) From 2 to 3

OpenStudy (campbell_st):

you need to find the ordered pairs find f(-5) and f(-3) so the average rate of change is \[\frac{f(-5) - f(-3)}{-5 - (-3)}\] you are really just finding the slope of the line between the 2 points

OpenStudy (anonymous):

Okay thanks!!

OpenStudy (anonymous):

okay that wasn't the right answer, but thanks!!!

OpenStudy (whpalmer4):

what wasn't the right answer?

OpenStudy (anonymous):

I just don't know how to do the problem that's all, getting frustrated.... Thanks!!!

OpenStudy (whpalmer4):

well, @campbell_st did tell you what to do. I'll try explaining it again. To find the average rate of change between two points on the curve, you need to find the (x,y) values for the two points. In the first part, they are (-5,f(-5)) and (-3,f(-3)). Then you plug those values into the formula for the slope of a line passing through the two points: \[m = \frac{y_2-y_1}{x_2-x_1}\]and \(m\) is your average rate of change. For the first part, that would be \[m = \frac{f(-5) - f(-3)}{-5-(-3)} = \frac{ ((-5)^3-3(-5)+1)) - ((-3)^3-3(-3)+1)) }{-2} \]\[=\frac{(-125+15+1)-(-27+9+1)}{-2} =\frac{ -109-(-17)}{-2} = \frac{-92}{-2} =46\]

OpenStudy (whpalmer4):

Now you repeat the drill for the other two sets of points. If you post your answers, I'll check them for you.

OpenStudy (anonymous):

okay......

OpenStudy (anonymous):

Never mind whpalmer, I'm just not getting it.. when it comes to math, you need a blended environment, meaning a classroom setting... I'm more of a visual learner, and math is my worst subjects.... I thank you for your time, that you have spent with me... Don't worry about the rest of the problems... Thanx!!!

OpenStudy (whpalmer4):

I'd hate to have you be the first person I couldn't successfully teach.

OpenStudy (whpalmer4):

let's break it down further. do you understand the notion of slope?

OpenStudy (whpalmer4):

(my previous remark was not intended as any sort of slight or judgment — just that I'm willing to work with anyone who wants to learn for as long as it takes)

OpenStudy (anonymous):

No you did a very good job trust me, I'm frustrated right now to where I can't think.. Math is a type of subject where you can't spend a long time on, especially on the internet. I'm kind of wore out, that's all... Trust me you did a fine job!!

OpenStudy (whpalmer4):

feel free to send me a message later if you want help with this or anything else.

OpenStudy (whpalmer4):

go outside, take a walk, ride your bike, go for a swim, do something else for a while. breaks are good :-)

OpenStudy (anonymous):

okay I will do that... Let me get a bite to eat.. I think that's whats wrong with me:)

OpenStudy (anonymous):

Thanks for the advice...... c u soon....